# KiDS-Legacy: Redshift distributions and their calibration

### v1

@@ 5.4 Modelling of the measurements
A long-standing issue of clustering redshifts is that the measured correlation functions are, by definition, no probability densities and must therefore be modelled in some way. A number of different approaches (e.g. Johnson et al., 2017; Stölzner et al., 2021; Gatti et al., 2022; Naidoo et al., 2023) have been implemented to mitigate the frequently arising negative correlation amplitudes, as they represent only a noisy realisation of the underlying redshift distribution. Given the dominant sensitivity of cosmic shear to the mean redshifts of the tomographic bins, we opt for simplicity here and use the SOM $N(z)$ to fit a simple shift $D_{\rm z}$ in redshift and a free normalisation121212 The normalisation is necessary because we cannot expect the data points to be properly normalised a priori due to galaxy bias or other systematic effects. $A$ such that
@@
This approach, however, can be quite sensitive to single data points with small variance, which can be aggravated in case of mismatches between the shape of the CCs and the model $N(z)$. This is of particular importance since there seems to be some additional intrinsic scatter in the data CCs that exceeds the variance that one expects given the uncertainties of the measurements. This intrinsic scatter is most obvious in bins five and six and in particular at high redshifts.
Observational systematics might introduce such erroneous additional correlations and could be reduced in future work by using organised random catalogues (Yan et al., 2024). Here, we opt for a simpler empirical correction and extend our fit model with an additive error term $f\,(1+z)$ with free amplitude $f$ such that the combined uncertainty $s$ for each measurement is
@@
We integrate this error term into our likelihood and marginalise over $f$ when determining the shift parameters $D_{\rm z}$ and the amplitudes $A$.
This process of measuring and modelling clustering redshifts with a reference $N(z)$ is tested on the MICE2 mocks, where we can additionally use the true redshift distribution as fit model in Eq. (5). This allows us to verify the robustness of and determine any biases inherent to our methodology. A key difference with respect to the data is however that the additional intrinsic scatter, parameterised by the $f$-term, is not present in MICE2. We therefore determine $f$ on the data by fitting each tomographic bin and add the expected intrinsic scatter to the mock measurements. We perturb the data points (but not their uncertainties, which were already adapted via the method described in Sect. 5.3) with Gaussian noise with a variance of $f^{2}\,(1+z)^{2}$, which is also included in the example shown in Fig. 8. Since we must avoid biases that may arise from a certain random realisation of the added scatter, we always create 100 realisations, fit each of them independently and compute the mean and variance of $D_{\rm z}$ from all realisations.
@@ 6 Simulation Results
In this section, we present the results of the redshift distribution estimation from our simulated galaxy samples. In Sect. 6.1 we detail the redshift distributions and bias parameters estimated using our SOM algorithm. In Sect. 6.2 the redshift distributions and bias parameters estimated using the CC method are shown. Within these sections, we cover the sensitivity tests that are performed to determine the robustness of the methods.
Table 5: Summary of redshift distribution estimates from simulations.
@@
$\langle\delta_{z}\rangle$0.026 0.014 -0.002 0.008 -0.011 -0.054
$\sigma_{\delta z}$ 0.002 0.001 0.002 0.002 0.002 0.004
@@
$\langle\delta_{z}\rangle$0.025 0.013 -0.001 0.008 -0.012 -0.053
@@
$\left[{\textbf{D}}\right]:$ No stellar contam. $\langle\hat{\mu}_{z}\rangle$ 0.316 0.470 0.608 0.800 1.008 1.334 6.1.4
@@
$\langle\delta_{z}\rangle$0.023 0.008 -0.001 0.006 -0.007 -0.039
@@
$\left[{\textbf{E}}\right]:$ No calib. weights $\langle\hat{\mu}_{z}\rangle$ 0.346 0.497 0.637 0.803 0.982 1.302 6.1.5
$\langle\delta_{z,0}\rangle$0.004 0.035 0.006 -$0$0.087 -0.163
@@
$\left[{\textbf{G}}\right]:$ Swgt only (no prior volume weight) $\langle\hat{\mu}_{z}\rangle$ 0.345 0.494 0.636 0.807 0.997 1.298 6.1.6
$\langle\delta_{z,0}\rangle$0.018 0.030 -0.002 -0.006 -0.096 -0.193
@@
$\left[{\textbf{F}}\right]:$ Pvol only (No calib. shape weight) $\langle\hat{\mu}_{z}\rangle$ 0.318 0.472 0.608 0.801 1.008 1.335 6.1.7
@@
$\langle\delta_{z}\rangle$0.020 0.013 0.004 0.011 -0.001 -0.032
@@
SKiLLS trunc. $\left[{\textbf{H}}\right]:$ Pvol only (no calib. shape weight) $\langle\hat{\mu}_{z}\rangle$ 0.320 0.469 0.604 0.798 0.983 1.110 6.1.8
@@
MICE2 $\left[{\textbf{I}}\right]:$ Pvol only (no calib. shape weight) $\langle\hat{\mu}_{z}\rangle$ 0.304 0.468 0.593 0.796 0.965 1.097 6.1.8
@@
131313 Entries above include abbreviations for prior volume weights (‘Pvol’), lensfit shape weights (‘Swgt’), Statistics in the table include: $\langle\hat{\mu}_{z}\rangle$: average of mean redshifts of all realisations of the calibration sample $\langle\delta_{z,0}\rangle$: average bias of the mean redshift prior to SOM $\langle\delta_{z}\rangle$: average bias of the mean redshift after SOM weighting $\sigma_{\delta z}$: uncertainty of $\langle\delta_{z}\rangle$, where values of $\leq 0.01$ are considered negligible. $\Delta\langle\delta_{z}\rangle$: difference in biases (after SOM weighting) between scenario and reference, i.e. $\langle\delta_{z}\rangle-\langle\delta_{z}\rangle_{\rm ref}$
@@ 6.1 SOM redshift distributions
Table 5 summarises the results of our SOM redshift distribution calibration using our various simulations. The table presents multiple statistics, which we outline here. $\langle\hat{\mu}_{z}\rangle$ is the average of the estimated redshift distribution first moments (i.e. means) under realisations of the calibration sample (see Sect. 3). $\langle\delta_{z,0}\rangle$ is the average bias in the first moments prior to any reweighting by the SOM, under realisations of the calibration sample, relative to the true weighted redshift distribution of the lensing sample (which is naturally unknowable for real data). $\langle\delta_{z}\rangle$ is the average bias in the first moments after reweighting by the SOM. $\sigma_{\delta z}$ is the standard deviation of the distribution of biases after reweighting by the SOM. Finally, $\Delta\langle\delta_{z}\rangle=\langle\delta_{z}\rangle-\langle\delta_{z}\rangle_{\rm ref}$ is the difference of the average biases, after reweighting by the SOM, between a given scenario and a reference/fiducial scenario.
@@ 6.1.1 $\left[{\textbf{A}}\right]:$ Fiducial calibration
We compute the $N(z)$ and bias parameters for our fiducial SOM redshift calibration methodology and parameter set (Table 4), and subsequently compare alternative analyses (such as sensitivity tests) to these fiducial results. The results of these sensitivity tests are then folded into our analysis in one of two ways. For tests that are expected to have no impact on the redshift distributions (such as perturbations to ad-hoc parameters of the simulation construction), we fold differences in recovered $N(z)$ into our systematic error budget. For tests that are expected to have some systematic (possibly non-negligible) impact on the redshift distributions (such as alternate sample weighting), we utilise the resulting $N(z)$ and bias parameters for use in full end-to-end reruns of the KiDS-Legacy cosmology, and present these results as alternative cosmological constraints in Wright et al. (submitted). This is because the differences in biases that arise from the latter type of analyses are not indicative of systematic effects in the methodology: rather the samples underlying the analysis have become systematically different.
Figure 9: Redshift distributions for the fiducial SOM calibration methodology, computed using the SKiLLS simulations. The ensemble of blue polygons show the $N(z)$ constructed with different realisations of the calibration sample. Redshift distributions include the new gold-weighting for the wide-field sample, and prior volume weighting and shape measurement weights for the calibration samples.
Resulting $N(z)$ for our fiducial SOM redshift calibration methodology are provided in Fig. 9. The redshift distributions are well constrained to the tomographic bin limits for all six tomographic bins used in KiDS-Legacy, with the third tomographic bin showing the largest “outlier population” (an extended tail to low redshifts) and the sixth tomographic bin showing a bias to somewhat lower redshifts than targeted by the photo-$z$ cuts. The figure shows the full range of redshift distributions estimated with realisations of the calibration sample as a filled polygon. Overlaid is the true target $N(z)$, which is also shown as a polygon but which has, in effect, vanishingly small area. The figure demonstrates the consistency of the estimated and true redshift distributions: in every bin, our estimated $N(z)$ are able to successfully capture the full complexity of the source redshift distributions with accuracy and precision.
@@
Finally, we note again that, compared to previous SOM redshift calibration work within KiDS (see, e.g., Wright et al., 2020a; Hildebrandt et al., 2021; van den Busch et al., 2022), the redshift calibration process here utilises different tomographic binning and a much larger spectroscopic calibration sample. Furthermore, we implement various weights and selections on the calibration side not previously implemented in KiDS. This makes direct comparison between our fiducial results and results presented in previous KiDS work difficult; it would be inappropriate, for example, to apply calibrations presented here to previous work with KiDS (without redoing the redshift calibration entirely).
@@ 6.1.2 $\left[{\textbf{B}}\right]:$ Single shear realisation
Our fiducial analysis uses the full SKiLLS simulation set, including eight realisations of the simulated photometry catalogues that are generated by applying four uniform shears and two position angle rotations to all sources in the $108$ $\text{deg}^{2}$ of simulated sky. These realisations are useful as each has an independent realisation of photometric noise, allowing us to increase the effective size of our simulated catalogue by a factor of eight. This does not, however, reduce the sample variance contribution to the analysis, as all photometric realisations are drawn from the same underlying large-scale structures.
The motivation for the use of all shear realisations in the fiducial pipeline is a practical one: we require the shear realisations for shape calibration, and all sources must be appropriately gold-weighted. However, using the multiple photometric realisations considerably increases the runtime of our calibration. Therefore, in the interest of speed and reducing unnecessary power consumption, we only utilise a single shear realisation for much of the redshift calibration testing presented here.
As such, we first verify that the results that we find for our fiducial redshift calibration process are unchanged under reduction of the number of sources available for calibration testing. These results are presented in Table 5: we find that the use of a single shear realisation produces bias estimates that are consistent with the fiducial case to better than $|\Delta\langle{\delta z}\rangle|=0.0012$ in all bins. Furthermore, the scatter of the calibrated biases is always consistent between the single and fiducial shear runs, with the exceptions that the scatter is increased in bin one and decreased in bin six. However, as the absolute size of the scatter is still relatively small, we conclude that there is unlikely to be a systematic bias introduced in our testing process by using a single shear realisation in redshift calibration testing.
@@ 6.1.3 $\left[{\textbf{C}}\right]:$ Matching algorithm validation
We validate the robustness of our calibration to the ad-hoc matching algorithm parameters (described in Sect. 3.3.1) by constructing multiple realisations of the calibration sample using perturbations on the fiducial choices. We then propagate these modified calibration samples through the redshift calibration pipeline and compare the biases that we estimate to those from the fiducial setup. For perturbations in the matching parameters, we test three choices of redshift window size ($1000$, $2000$, and 10 0001000010\,00010 000 sources respectively), 10 perturbations to the matching algorithm feature space (dropping one band at a time), matching on colours and a reference magnitude, and matching without photo-$z$ information. We find that the matching algorithm is extremely robust to each of these perturbations, with typical changes in the bias (from that measured in the fiducial case) of order $\Delta\langle{\delta z}\rangle=10^{-4}$. For use in our uncertainty budget, we compute the standard deviation of the recovered bias parameters under the different realisations of the matching algorithm; these are reported in Table 5. In all bins the scatter introduced in the bias from the matching algorithm perturbations is less than half the scatter in the spatial/photometric realisations of the calibrations samples in the fiducial case (i.e. $\sigma_{\Delta\langle{\delta z}\rangle}<0.5\sigma_{\langle{\delta z}\rangle,{\rm fid}}$). The maximal uncertainty in the bias introduced by our matching algorithm perturbations is $\sigma_{\Delta\langle{\delta z}\rangle}=0.0029$, in the sixth tomographic bin.
6.1.4 $\left[{\textbf{D}}\right]:$ Impact of stellar contamination

### v2

@@ 5.4 Modelling of the measurements
A long-standing issue within clustering redshift measurements is that the measured correlation functions are, by definition, not probability densities and must therefore be modelled in some way. A number of different approaches (e.g. Johnson et al., 2017; Stölzner et al., 2021; Gatti et al., 2022; Naidoo et al., 2023) have been implemented to mitigate the frequently arising negative correlation amplitudes, as they represent only a noisy realisation of the underlying redshift distribution. Given the dominant sensitivity of cosmic shear to the mean redshifts of the tomographic bins, we opted for simplicity here and use the SOM $N(z)$ as a model that we fit to the CC data via a simple shift in redshift, $D_{\rm z}$, and a free normalisation [^12] The normalisation is necessary because we cannot expect the data points to be properly normalised a priori due to galaxy bias or other systematic effects. , $A$, such that
@@
This approach, however, can be quite sensitive to single data points with small variance, which can be aggravated in case of mismatches between the shape of the CCs and the model $N(z)$. This is of particular importance since there seems to be some additional intrinsic scatter in the data CCs that exceeds the variance that one expects given the uncertainties of the measurements. This intrinsic scatter is most obvious in bins five and six and, in particular, at high redshifts.
Observational systematics might introduce such erroneous additional correlations and could be reduced in future work by using organised random catalogues (Yan et al., 2025). Here, we opted for a simpler empirical correction and extended our fit model with an additive error term f(1+z)f\,(1+z) with free amplitude, $f$, such that the combined uncertainty $s$ for each measurement was
@@
We integrated this error term into our likelihood and marginalised over $f$ when determining the shift parameters $D_{\rm z}$ and the amplitudes $A$.
This process of measuring and modelling clustering redshifts with a reference $N(z)$ was tested on the MICE2 mocks, where we could additionally use the true redshift distribution as a fit model in Eq. (5). This allowed us to verify the robustness of, and determine any biases inherent to, our methodology. A key difference with respect to the data, however, is that the additional intrinsic scatter, parameterised by the $f$-term, is not present in MICE2. We therefore determined $f$ on the data by fitting each tomographic bin and added the expected intrinsic scatter to the mock measurements. We perturbed the data points (but not their uncertainties, which were already adapted via the method described in Sect. 5.3) with Gaussian noise with a variance of $f^{2}\,(1+z)^{2}$, which is also included in the example shown in Fig. 8. Since we must avoid biases that may arise from a certain random realisation of the added scatter, we always created 100 realisations, fit each of them independently, and computed the mean and variance of $D_{\rm z}$ from all realisations.
@@ 6 Simulation results
In this section, we present the results of the redshift distribution estimation from our simulated galaxy samples. In Sect. 6.1 we detail the redshift distributions and bias parameters estimated using our SOM algorithm. In Sect. 6.2 the redshift distributions and bias parameters estimated using the CC method are shown. Within these sections, we cover the sensitivity tests that were performed to determine the robustness of the methods.
Table 5: Redshift distribution estimates from simulations. [^13] Entries above include abbreviations for prior volume weights (‘Pvol’), lensfit shape weights (‘Swgt’), Statistics in the table include: $\langle\hat{\mu}_{z}\rangle$: average of mean redshifts of all realisations of the calibration sample $\langle\delta_{z,0}\rangle$: average bias of the mean redshift prior to SOM $\langle\delta_{z}\rangle$: average bias of the mean redshift after SOM weighting $\sigma_{\delta z}$: uncertainty of $\langle\delta_{z}\rangle$, where values of $\leq 0.01$ are considered negligible. $\Delta\langle\delta_{z}\rangle$: difference in biases (after SOM weighting) between scenario and reference, i.e. $\langle\delta_{z}\rangle-\langle\delta_{z}\rangle_{\rm ref}$
@@
$\langle\delta_{z}\rangle$ -0.026 0.014 -0.002 0.008 -0.011 -0.054
$\sigma_{\delta z}$ 0.002 0.001 0.002 0.001 0.002 0.004
@@
$\langle\delta_{z}\rangle$ -0.025 0.013 -0.001 0.008 -0.012 -0.053
@@
$\left[{\textbf{D}}\right]:$ $z$-dependant Shear $\langle\hat{\mu}_{z}\rangle$ 0.314 0.470 0.603 0.798 0.999 1.313 6.1.4
$\langle\delta_{z}\rangle$ -0.027 0.012 -0.003 0.008 -0.010 -0.051
$\sigma_{\delta z}$ 0.002 0.001 0.002 0.001 0.002 0.005
$\Delta\langle\delta_{z}\rangle$ 0.001 0.002 0.001 0.000 -0.001 -0.003
$\left[{\textbf{E}}\right]:$ No stellar contam. $\langle\hat{\mu}_{z}\rangle$ 0.316 0.470 0.608 0.800 1.008 1.334 6.1.5
@@
$\langle\delta_{z}\rangle$ -0.023 0.008 -0.001 0.006 -0.007 -0.039
@@
$\left[{\textbf{F}}\right]:$ No calib. weights $\langle\hat{\mu}_{z}\rangle$ 0.346 0.497 0.637 0.803 0.982 1.302 6.1.6
$\langle\delta_{z,0}\rangle$ -0.004 0.035 0.006 -0.000 -0.087 -0.163
@@
$\left[{\textbf{G}}\right]:$ Swgt only (no prior volume weight) $\langle\hat{\mu}_{z}\rangle$ 0.345 0.494 0.636 0.807 0.997 1.298 6.1.7
$\langle\delta_{z,0}\rangle$ -0.018 0.030 -0.002 -0.006 -0.096 -0.193
@@
$\left[{\textbf{H}}\right]:$ Pvol only (No calib. shape weight) $\langle\hat{\mu}_{z}\rangle$ 0.318 0.472 0.608 0.801 1.008 1.335 6.1.8
@@
$\langle\delta_{z}\rangle$ -0.020 0.013 0.004 0.011 -0.001 -0.032
@@
SKiLLS trunc. $\left[{\textbf{I}}\right]:$ Pvol only (no calib. shape weight) $\langle\hat{\mu}_{z}\rangle$ 0.320 0.469 0.604 0.798 0.983 1.110 6.1.9
@@
MICE2 $\left[{\textbf{J}}\right]:$ Pvol only (no calib. shape weight) $\langle\hat{\mu}_{z}\rangle$ 0.304 0.468 0.593 0.796 0.965 1.097 6.1.9
@@ 6.1 SOM redshift distributions
Table 5 summarises the results of our SOM redshift distribution calibration using our various simulations. The table presents multiple statistics, which we outline here. $\langle\hat{\mu}_{z}\rangle$ is the average of the estimated redshift distribution first moments (i.e. means) under realisations of the calibration sample (see Sect. 3). $\langle\delta_{z,0}\rangle$ is the average bias in the first moments prior to any re-weighting by the SOM, under realisations of the calibration sample, relative to the true weighted redshift distribution of the lensing sample (which is naturally unknowable for real data). $\langle\delta_{z}\rangle$ is the average bias in the first moments after re-weighting by the SOM. $\sigma_{\delta z}$ is the standard deviation of the distribution of biases after re-weighting by the SOM. Finally, $\Delta\langle\delta_{z}\rangle=\langle\delta_{z}\rangle-\langle\delta_{z}\rangle_{\rm ref}$ is the difference of the average biases, after re-weighting by the SOM, between a given scenario and a reference/fiducial scenario.
@@ 6.1.1 $\left[{\textbf{A}}\right]:$ Fiducial calibration
We computed the $N(z)$ and bias parameters for our fiducial SOM redshift calibration methodology and parameter set (Table 4), and subsequently compared alternative analyses (such as sensitivity tests) to these fiducial results. The results of these sensitivity tests were then folded into our analysis in one of two ways. For tests that are expected to have no impact on the redshift distributions (such as perturbations to ad hoc parameters of the simulation construction), we folded differences in recovered $N(z)$ into our systematic error budget. For tests that are expected to have some systematic (possibly non-negligible) impact on the redshift distributions (such as alternate sample weighting), we utilised the resulting $N(z)$ and bias parameters for use in full end-to-end reruns of the KiDS-Legacy cosmology, and presented these results as alternative cosmological constraints in Wright et al. (2025). This is because the differences in biases that arise from the latter type of analyses are not indicative of systematic effects in the methodology: rather the samples underlying the analyses have become systematically different.
Figure 9: Redshift distributions for the fiducial SOM calibration methodology, computed using the SKiLLS simulations. The ensemble of blue polygons show the $N(z)$ constructed with different realisations of the calibration sample. Redshift distributions include the new gold-weighting for the wide-field sample, and prior volume weighting and shape-measurement weights for the calibration samples.
Resulting $N(z)$ for our fiducial SOM redshift calibration methodology are provided in Fig. 9. The redshift distributions are well constrained to the tomographic bin limits for all six tomographic bins used in KiDS-Legacy, with the third tomographic bin showing the largest outlier population (an extended tail to low redshifts) and the sixth tomographic bin showing a bias to somewhat lower redshifts than targeted by the photo-$z$ cuts. The figure shows the full range of redshift distributions estimated with realisations of the calibration sample as a filled polygon. Overlaid is the true target $N(z)$, which is also shown as a polygon but which has, in effect, vanishingly small area. The figure demonstrates the consistency of the estimated and true redshift distributions: in every bin, our estimated $N(z)$ are able to successfully capture the full complexity of the source redshift distributions with accuracy and precision.
@@
Finally, we note again that, compared to previous SOM redshift calibration work within KiDS (see e.g. Wright et al., 2020a; Hildebrandt et al., 2021; van den Busch et al., 2022), the redshift calibration process here utilises different tomographic binning and a much larger spectroscopic calibration sample. Furthermore, we implement various weights and selections on the calibration side not previously implemented in KiDS. This makes direct comparison between our fiducial results and results presented in previous KiDS work difficult; it would be inappropriate, for example, to apply calibrations presented here to previous work with KiDS (without redoing the redshift calibration entirely).
@@ 6.1.2 $\left[{\textbf{B}}\right]:$ Single shear realisation
Our fiducial analysis used the full SKiLLS simulation set, including eight realisations of the simulated photometry catalogues that were generated by applying four uniform shears and two position angle rotations to all sources in the $108$ $\text{deg}^{2}$ of simulated sky. These realisations are useful as each has an independent realisation of photometric noise, allowing us to increase the effective size of our simulated catalogue by a factor of 8. This does not, however, reduce the sample variance contribution to the analysis, as all photometric realisations are drawn from the same underlying large-scale structures.
The motivation for the use of all shear realisations in the fiducial pipeline is a practical one: we require the shear realisations for shape calibration, and all sources must be appropriately gold-weighted. However, using the multiple photometric realisations considerably increases the runtime of our calibration. Therefore, in the interest of speed and reducing unnecessary power consumption, we only utilised a single shear realisation for much of the redshift calibration testing presented here.
In that respect, we first verified that the results that we find for our fiducial redshift calibration process are unchanged under reduction of the number of sources available for calibration testing. These results are presented in Table 5: we found that the use of a single shear realisation produces bias estimates that are consistent with the fiducial case to better than $|\Delta\langle{\delta z}\rangle|=0.0012$ in all bins. Furthermore, the scatter of the calibrated biases is always consistent between the single and fiducial shear runs, with the exceptions that the scatter is increased in bin one and decreased in bin six. However, as the absolute size of the scatter is still relatively small, we conclude that there is unlikely to be a systematic bias introduced in our testing process by using a single shear realisation in redshift calibration testing.
@@ 6.1.3 $\left[{\textbf{C}}\right]:$ Matching algorithm validation
We validated the robustness of our calibration to the ad hoc matching algorithm parameters (described in Sect. 3.3.1) by constructing multiple realisations of the calibration sample using perturbations on the fiducial choices. We then propagated these modified calibration samples through the redshift calibration pipeline and compared the biases that we estimated to those from the fiducial setup. For perturbations in the matching parameters, we tested three choices of redshift window size ($1000$, $2000$, and 10 00010\,000 sources, respectively), 10 perturbations to the matching algorithm feature space (dropping one band at a time), matching on colours and a reference magnitude, and matching without photo-$z$ information. We found that the matching algorithm is extremely robust to each of these perturbations, with typical changes in the bias (from that measured in the fiducial case) of order $\Delta\langle{\delta z}\rangle=10^{-4}$. For use in our uncertainty budget, we computed the standard deviation of the recovered bias parameters under the different realisations of the matching algorithm; these are reported in Table 5. In all bins the scatter introduced in the bias from the matching algorithm perturbations is less than half the scatter in the spatial/photometric realisations of the calibrations samples in the fiducial case (i.e. $\sigma_{\Delta\langle{\delta z}\rangle}<0.5\sigma_{\langle{\delta z}\rangle,{\rm fid}}$). The maximal uncertainty in the bias introduced by our matching algorithm perturbations is $\sigma_{\Delta\langle{\delta z}\rangle}=0.0029$, in the sixth tomographic bin.
6.1.4 $\left[{\textbf{D}}\right]:$ Impact of redshift-dependent shear
Recent analyses by MacCrann et al. (2022) highlighted the interplay between redshift distribution estimation and shape-measurement calibration for ongoing stage-III cosmic shear surveys. This was a primary motivating factor for the construction of the SKiLLS simulations by Li et al. (2023), who used the redshift-dependent shear realisations of SKiLLS to estimate the impact of higher-order blending effects on an analysis such as KiDS-Legacy. The found a measurable, but otherwise minor, impact of redshift-dependent shear and blending on the computation of shape-measurement bias in KiDS-Legacy. We repeated this analysis in the context of redshift distribution bias estimation.
In the context of redshift distribution estimation, redshift-dependent shear primarily manifests as a modification to the weights of sources, which are blended across significant redshift baselines. This is therefore a fairly minor perturbation to the properties used to estimate our redshift distributions and, as a result, the impact that this effect has on our recovered $N(z)$ is similarly minor. In all bins, the measured change in bias, which we find from analysing SKiLLS simulations with redshift-dependent shear, is $\Delta\langle\delta_{z}\rangle<0.003$. Therefore, we conclude that redshift-dependent shear in blends is a subdominant source of bias, for both shape-calibration and redshift-distribution calibration, at the sensitivity of KiDS-Legacy.
6.1.5 $\left[{\textbf{E}}\right]:$ Impact of stellar contamination